OpenStudy (anonymous):

Find the factored form of this trinomial. a^2-10a-24

zepdrix (zepdrix):

This will factor into two binomials. \[\large a^2-10a-24\] We want `factors` of -24 that will `add` to -10. Here are a couple factors we can try, \[\large -6 \times 4 =-24\]But when we try to add these numbers,\[\large -6+4\neq -10\] They didn't `add` up to -10. Hmm those didn't work out for us. We need to try some other factors.

zepdrix (zepdrix):

How about these maybe.\[\large -12 \times 2\]or\[\large 12 \times (-2)\] Maybe one of these sets of factors will work :D Hmmm

zepdrix (zepdrix):

\[\large -12 \times 2 = -24\]Adding them we get,\[\large \color{orangered}{-12}+\color{royalblue}{2}=-10\]Yay it worked! So we have determined that these factors will be included in our binomials. \[\large a^2-10a-24 \qquad = \qquad (a\color{orangered}{-12})(a+\color{royalblue}{2})\]

OpenStudy (anonymous):

ty